# Reconstruction of strong degeneracy region for parabolic equations and systems

### v1

@@ 8 Numerical results
In this section, we will show some numerical results related to the identification of the degeneracy point $a\in(0,1)$ and also the initial data $(u_{0},v_{0})$ in
@@
We will perform some numerical tests for the strong degeneracy case with $\theta\in[1,2)$, which have not been considered in the previous article [5], where only the scalar case with $\theta=1$ has been treated.
In particular, this Section will be devoted to the numerical reconstruction of the solution for several kinds of inverse problem. More precisely, the following two tests will consider the inverse problem of recovering the degeneracy point, for both one-point and distributed measurements.
@@
The other tests will consider the more general inverse problem of degeneracy and initial data reconstruction, taking into account distributed measurements on one or two sides of the domain. These tests, especially Test $4$ and Test $6$, also illustrate the uniqueness theoretical results of Section 6 and Section 7.
@@
Find $a$ and constant initial data $(u_{0},v_{0})$ from the distributed measurements $\eta(t)=\partial_{x}u(1,t)$ and $\zeta(t)=\partial_{x}v(1,t)$ for $t\in(t_{1},t_{2})$.
@@ 8.1 Degeneracy reconstruction with one-point measurements
and a functional $J:a\in\mathcal{U}^{a}_{ad}\mapsto\mathds{R}$ given by
$J(a)=\dfrac{1}{2}\displaystyle|\eta(t^{\ast})-\partial_{x}u^{a}(1,t^{\ast})|^{2}+\dfrac{1}{2}\displaystyle|\zeta(t^{\ast})-\partial_{x}v^{a}(1,t^{\ast})|^{2}$
for some $t^{\ast}\in(0,T)$. The related optimization problem is the following:
@@
The goal is to reconstruct the degeneracy point $a$ for a strong degenerate case and with the given initial data $u_{0}=1$, $v_{0}=1$.
We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=4$, $t^{\ast}=1.99$ and $aini=0.1$ as initial guess to recover the desired value of $a_{d}=0.5$ by the minimization algorithm.
The numerical results can be seen in Figures 5 and 5. The round points correspond to iterations during the optimization algorithm and the digits show the number of iterations performed. With the solid line, we have represented the evolution of the cost. We obtain the computed value $a_{c}=0.4999999999999895$ and the cost $J(a_{c})\approx 1.e-27$.
@@
Figures 7 and 7 show the results obtained for $\theta=1.01$ and $aini=0.1$. We can see that for smaller values of $\theta$, we need more iterations to achieve the convergence. The computed value is $a_{c}=5000270641466984$ and the cost $J(a_{c})\approx 1.e-11$ obtained in iteration 18.
Figure 6: Test 1, $\theta=1.01$, $t^{\ast}=1.99$, $u_{0}=1$, $v_{0}=1$. Iterations in the computation of $a$ by trust-region-reflective algorithm, $aini=0.1$.
Figure 7: Test 1, $\theta=1.01$, $t^{\ast}=1.99$, $u_{0}=1$, $v_{0}=1$. Evolution of the cost in trust-region-reflective algorithm, $aini=0.1$.
@@ 8.2 Degeneracy reconstruction with distributed measurements
where $\mathcal{U}^{a}_{ad}$ is given by (48) and $\mathcal{I}:a\in\mathcal{U}^{a}_{ad}\mapsto\mathds{R}$ is defined as follows
$\mathcal{I}(a)=\dfrac{1}{2}\displaystyle\int_{t_{1}}^{t_{2}}|\eta(t)-\partial_{x}u^{a}(1,t)|^{2}\,dt.$
@@
We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=4$, $t_{1}=0$, $t_{2}=T$, $u_{0}=1$, $v_{0}=1$ and $aini=0.1$ as initial guess to recover the desired value of $a_{d}=0.5$ using the minimization algorithm.
The numerical results can be seen in Figures 9 and 9. The round points correspond to iterations during the optimization algorithm. With the solid line, we have represented the evolution of the cost. We obtain the computed value $a_{c}=0.4999999999999999$ and the cost $\mathcal{I}(a_{c})\approx 1.e-26$. We can see in Table 2 that this procedure is also stable with respect to random perturbations in the target.
@@
Noise Cost Iterations $a_{c}$
1% 1.e-16 17 0.4994683282640736
0.1% 1.e-13 11 0.4999673889973712
0.01% 1.e-11 7 0.5000095945877255
0% 1.e-26 11 0.4999999999999999
Table 2: Evolution of the cost with random noises in the target, Test 2 with $\theta=1.5$ and $aini=0.1$.
@@ 8.3 Degeneracy and initial data reconstruction with distributed measurements
In this section, we will give some numerical simulations of reconstruction of the degeneracy point $a\in(0,1)$ and initial data $(u_{0},v_{0})$ by distributed measurements. The following analysis will allow us to better highlight the difference between the two uniqueness Theorems in Section 6 (or Section 7 for the system version), in terms of the results for the initial data.
8.3.1 Constant initial data, one side measurements
Let us assume that the initial data $(u_{0},v_{0})$ are constant on the whole interval. We will see that, in this case, we can perform the reconstruction using the distributed measurements only on one side of the interval $(0,1)$. Therefore, our goal is to find $a\in(0,1)$ and initial data $(u_{0},v_{0})$ such that a solution to (47) satisfies
[…]
where $\mathcal{U}^{a}_{ad}$ is given by (48) and $\mathcal{G}:(a,u_{0},v_{0})\in\mathcal{U}^{a}_{ad}\times\mathds{R}\times\mathds{R}\mapsto\mathds{R}$ is defined as follows:
$\mathcal{G}(a,u_{0},v_{0})=\dfrac{1}{2}\displaystyle\int_{t_{1}}^{t_{2}}|\eta(t)-\partial_{x}u^{a,u_{0},v_{0}}(1,t)|^{2}\,dt+\dfrac{1}{2}\displaystyle\int_{t_{1}}^{t_{2}}|\zeta(t)-\partial_{x}v^{a,u_{0},v_{0}}(1,t)|^{2}\,dt.$
Test 3
We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=4$, $t_{1}=0$, $t_{2}=T$, $(u_{0ini},v_{0ini})=(0.5,1.5)$ and $aini=0.9$ as initial guesses to recover the desired value of $a_{d}=0.5$, $(u_{0d},v_{0d})=(1,2)$ using the minimization algorithm.
The numerical results can be seen in Figures 11,  11 and 12. In Figure 11, the round points again correspond to iterations during the optimization algorithm in the computation of $a$. In Figure 11, the stars and squares represent iterations during the optimization algorithm in the computation of $u_{0}$ and $v_{0}$. With the solid line in Figure 12, we have represented the evolution of the cost. We obtain the computed values $a_{c}=0.5000000000000001$, $u_{0c}=1.0000000000000011$, $v_{0c}=2.0000000000000062$ and the cost $\mathcal{G}(a_{c},u_{0c},v_{0c})\approx 1.e-25$.
Figure 10: Test 3, $\theta=1.5$. Iterations in the computation of $a$ by trust-region-reflective algorithm, $aini=0.9$, $u_{0ini}=0.5$, $v_{0ini}=1.5$.
Figure 11: Test 3, $\theta=1.5$. Iterations in the computation of $u_{0}$ and $v_{0}$ by trust-region-reflective algorithm, $aini=0.9$, $u_{0ini}=0.5$, $v_{0ini}=1.5$.
Figure 12: Test 3, $\theta=1.5$. Evolution of the cost in trust-region-reflective algorithm, $aini=0.9$, $u_{0ini}=0.5$, $v_{0ini}=1.5$.
@@ 8.3.2 Piecewise-constant initial data, two side measurements
Test 4
We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=4$, $t_{1}=0$, $t_{2}=T$, $u_{01ini}=0.5$, $u_{02ini}=1.5$, and $aini=0.1$ as initial guesses to recover the desired value of $a_{d}=0.5$, $u_{01d}=1$, $u_{02d}=2$ using the minimization algorithm.
The numerical results can be seen in Figures 14,  14 and 15. In Figure 14, the stars and the round points correspond to the iterations during the optimization algorithm in the computation of $u_{01}$ and $u_{02}$, respectively. In Figure 14, the round points represent the iterations during the optimization algorithm in the computation of $a$. With the solid line in Figure 15, we have represented the evolution of the cost during iterations. We obtain the computed value $a_{c}=0.5000000000049045$, $u_{01c}=1.0000000000141918$, $u_{02c}=1.9999999999341966$ and the cost $\mathcal{H}(a_{c},u_{01c},u_{02c})\approx 1.e-18$. This test allows us to numerically get the uniqueness result of Theorem 6.1.
Figure 13: Test 4, $\theta=1.5$. Iterations in the computation of $u_{01}$ and $u_{02}$ by trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.5$.
Figure 14: Test 4, $\theta=1.5$. Iterations in the computation of $a$ by trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.5$.
Figure 15: Test 4, $\theta=1.5$. Evolution of the cost in trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.5$.
@@ 8.3.3 Piecewise-constant initial data, one side measurements
We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=2$, $t_{1}=0$, $t_{2}=T$, $u_{01ini}=0.5$, $u_{02ini}=1.8$ and $aini=0.1$ as initial guesses to recover the desired value of $a_{d}=0.5$, $u_{01d}=1$, $u_{02d}=2$ using the minimization algorithm.
The numerical results can be seen in Figures 17,  17 and 18. In Figure 17, the stars and the round points correspond to the iterations during the optimization algorithm in the computation of $u_{01}$ and $u_{02}$, respectively. In Figure 17, the round points represent the iterations during the optimization algorithm in the computation of $a$. With the solid line in Figure 18, we have represented the evolution of the cost during iterations. The algorithm does not converge well and we can see that the value of $u_{01}$ is not recovering properly. The value of the functional does not become small: this indicates that, in this case, we are not able to obtain a solution of the inverse problem. However, for the value of $u_{02}$ defined on the side where we take the distributed measurements, we get a better approximation. This suggests that the inverse problem might be solved in the right interval $(a,1)$, as we expect from the uniqueness Theorems 6.2 and 7.3.
Figure 16: Test 5, $\theta=1.5$. Iterations in the computation of $u_{01}$ and $u_{02}$ by trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.8$.
Figure 17: Test 5, $\theta=1.5$. Iterations in the computation of $a$ in trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.8$.
Figure 18: Test 5, $\theta=1.5$. Evolution of the cost in trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.8$.
In the next test, we will numerically obtain a result in line with the theoretical results 6.2 and 7.3. Hence, we fix the initial data $u_{01}$ in $(0,a)$ and leave $a$ and the initial datum $u_{02}$ in $(a,1)$ as unknown, reconstructing them based on distributed measurements of the normal derivative at $x=1$.
The reformulation of the inverse problem is now as follows:
@@
where $\mathcal{U}^{a}_{ad}$ is given by (48) and $\mathcal{K}:(a,u_{02})\in\mathcal{U}^{a}_{ad}\times\mathds{R}\times\mapsto\mathds{R}$ is defined as follows:
@@
We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=2$, $t_{1}=0$, $t_{2}=T$, $v_{0}=0$, $u_{01}=1$, $u_{02ini}=1.8$ and $aini=0.1$ as initial guesses to recover the desired value of $a_{d}=0.5$, $u_{02d}=2$ using the minimization algorithm.
The numerical results can be seen in Figures 20, 20 and 21. In Figures 20 and 20, the stars and round points correspond to iterations during the optimization algorithm in the computation of $u_{02}$ and $a$, respectively. With the solid line in Figure 21, we have represented the evolution of the cost during iterations. The algorithm converges well, in particular, we can see an appropriate reconstruction of the value $u_{02}$ defined on the side where we take the distributed measurements. We obtain the computed values $a_{c}=0.4999999996354925$, $u_{02c}=2.0000000002802012$ and the cost $\mathcal{K}(a_{c},u_{02c})\approx 1.e-18$.
Figure 19: Test 6, $\theta=1.5$. Iterations in the computation of $u_{02}$ by trust-region-reflective algorithm, $aini=0.1$, $u_{02ini}=1.8$.
Figure 20: Test 6, $\theta=1.5$. Iterations in the computation of $a$ in trust-region-reflective algorithm, $aini=0.1$, $u_{02ini}=1.8$.
Figure 21: Test 6, $\theta=1.5$. Evolution of the cost in trust-region-reflective algorithm, $aini=0.1$ and $u_{02ini}=1.8$.
@@
Acknowledgments
P. C. and V. D. were partially supported by the MUR Excellence Department Project awarded to the Department of Mathematics, University of Rome Tor Vergata, CUP E83C23000330006 and by the PRIN 2022 PNRR-Project P20225SP98 “Some mathematical approaches to climate change and its impacts” (funded by the European Community-Next Generation EU, CUP E53D2301791 0001). V. D. was also supported by the PRIN Project 2022FPZEES “Stability in Hamiltonian Dynamics and Beyond”. P. C. was also supported by the INdAM (Istituto Nazionale di Alta Matematica) Group for Mathematical Analysis, Probability and Applications. A. D. was partially supported by Grants PID2020-114976GB-I00 and PID2024-158206NB-I00, funded by MICIU/AEI (Spain). The authors warmly thank P. Martinez for enlightening discussions and his precious comments.

### v2

@@ 8 Numerical results
In this section, we will show some numerical results related to the identification of the degeneracy point $a\in(0,1)$, the initial data $(u_{0},v_{0})$ and also the degeneracy power $\theta$ in
@@
We will perform some numerical tests for the strong degeneracy case with $\theta\in[1,2)$, which have not been considered in the previous article [7], where only the scalar case with $\theta=1$ has been treated. In particular, this Section will be devoted to the numerical reconstruction of the solution for several kinds of inverse problem. More precisely, the following two tests will consider the inverse problem of recovering the degeneracy point, for both one-point and distributed measurements.
@@
The following three tests will consider the more general inverse problem of degeneracy and initial data reconstruction, taking into account distributed measurements on one or two sides of the domain. These tests, especially Test $3$ and Test $5$, also illustrate the first two uniqueness theoretical results of Section 6 and Section 7.
@@
The last test concerns the inverse problem of reconstructing the degeneracy power $\theta$ and the initial data from distributed measurements and covers the last uniqueness result in Section 6 and Section 7.
@@
Find $\theta$ and initial data $(u_{0},v_{0})$ in $(a,1)$ from the distributed measurements $\eta(t)=\partial_{x}u(1,t)$ and $\zeta(t)=\partial_{x}v(1,t)$ for $t\in(t_{1},t_{2})$.
@@ 8.1 Degeneracy reconstruction with one-point measurements
and a functional $J:a\in\mathcal{U}^{a}_{ad}\mapsto\mathds{R}$ given by $J(a)=\dfrac{1}{2}\displaystyle|\eta(t^{\ast})-\partial_{x}u^{a}(1,t^{\ast})|^{2}+\dfrac{1}{2}\displaystyle|\zeta(t^{\ast})-\partial_{x}v^{a}(1,t^{\ast})|^{2}$ for some $t^{\ast}\in(0,T)$. The related optimization problem is the following:
@@
The goal is to reconstruct the degeneracy point $a$ for a strong degenerate case and with the given initial data $u_{0}=1$, $v_{0}=1$. We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=4$, $t^{\ast}=1.99$ and $aini=0.1$ as initial guess to recover the desired value of $a_{d}=0.5$ by the minimization algorithm. The numerical results can be seen in Figures 5 and 5. The round points correspond to iterations during the optimization algorithm and the digits show the number of iterations performed. With the solid line, we have represented the evolution of the cost. We obtain the computed value $a_{c}=0.4999999999999895$ and the cost $J(a_{c})\approx 1.e-27$.
@@ 8.2 Degeneracy reconstruction with distributed measurements
where $\mathcal{U}^{a}_{ad}$ is given by (52) and $\mathcal{I}:a\in\mathcal{U}^{a}_{ad}\mapsto\mathds{R}$ is defined as $\mathcal{I}(a)=\dfrac{1}{2}\displaystyle\int_{t_{1}}^{t_{2}}|\eta(t)-\partial_{x}u^{a}(1,t)|^{2}\,dt$.
@@
We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=4$, $t_{1}=0$, $t_{2}=T$, $u_{0}=1$, $v_{0}=1$ and $aini=0.1$ as initial guess to recover the desired value of $a_{d}=0.5$ using the minimization algorithm. The numerical results can be seen in Figures 7 and 7. The round points correspond to iterations during the optimization algorithm. With the solid line, we have represented the evolution of the cost. We obtain the computed value $a_{c}=0.4999999999999999$ and the cost $\mathcal{I}(a_{c})\approx 1.e-26$.
@@ 8.3 Degeneracy and initial data reconstruction with distributed measurements
In this section, we will give some numerical simulations of reconstruction of the degeneracy point $a\in(0,1)$ and initial data $(u_{0},v_{0})$ by distributed measurements. The following analysis will allow us to better highlight the difference between the first two uniqueness Theorems in Section 6 (or Section 7 for the system version), in terms of the results for the initial data.
@@ 8.3.1 Piecewise-constant initial data, two side measurements
Test 3
We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=4$, $t_{1}=0$, $t_{2}=T$, $u_{01ini}=0.5$, $u_{02ini}=1.5$, and $aini=0.1$ as initial guesses to recover the desired value of $a_{d}=0.5$, $u_{01d}=1$, $u_{02d}=2$ using the minimization algorithm. The numerical results can be seen in Figures 9,  9 and 10. In Figure 9, the stars and the round points correspond to the iterations during the optimization algorithm in the computation of $u_{01}$ and $u_{02}$, respectively. In Figure 9, the round points represent the iterations during the optimization algorithm in the computation of $a$. With the solid line in Figure 10, we have represented the evolution of the cost during iterations. We obtain the computed value $a_{c}=0.5000000000049045$, $u_{01c}=1.0000000000141918$, $u_{02c}=1.9999999999341966$ and the cost $\mathcal{H}(a_{c},u_{01c},u_{02c})\approx 1.e-18$. This test allows us to numerically get the uniqueness result of Theorem 6.1.
Figure 8: Test 3, $\theta=1.5$. Iterations in the computation of $u_{01}$ and $u_{02}$ by trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.5$.
Figure 9: Test 3, $\theta=1.5$. Iterations in the computation of $a$ by trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.5$.
Figure 10: Test 3, $\theta=1.5$. Evolution of the cost in trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.5$.
@@ 8.3.2 Piecewise-constant initial data, one side measurements
Test 4
We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=2$, $t_{1}=0$, $t_{2}=T$, $u_{01ini}=0.5$, $u_{02ini}=1.8$ and $aini=0.1$ as initial guesses to recover the desired value of $a_{d}=0.5$, $u_{01d}=1$, $u_{02d}=2$ using the minimization algorithm. The numerical results can be seen in Figures 12,  12 and 13. In Figure 12, the stars and the round points correspond to the iterations during the optimization algorithm in the computation of $u_{01}$ and $u_{02}$, respectively. In Figure 12, the round points represent the iterations during the optimization algorithm in the computation of $a$. With the solid line in Figure 13, we have represented the evolution of the cost during iterations. The algorithm does not converge well and we can see that the value of $u_{01}$ is not recovering properly. The value of the functional does not become small: this indicates that, in this case, we are not able to obtain a solution of the inverse problem. However, for the value of $u_{02}$ defined on the side where we take the distributed measurements, we get a better approximation. This suggests that the inverse problem might be solved in the right interval $(a,1)$, as we expect from the uniqueness Theorems 6.2 and 7.3.
Figure 11: Test 4, $\theta=1.5$. Iterations in the computation of $u_{01}$ and $u_{02}$ by trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.8$.
Figure 12: Test 4, $\theta=1.5$. Iterations in the computation of $a$ in trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.8$.
Figure 13: Test 4, $\theta=1.5$. Evolution of the cost in trust-region-reflective algorithm, $aini=0.1$, $u_{01ini}=0.5$, $u_{02ini}=1.8$.
In the next test, we will numerically obtain a result in line with the theoretical results 6.2 and 7.3. Hence, we fix the initial data $u_{01}$ in $(0,a)$ and leave $a$ and the initial datum $u_{02}$ in $(a,1)$ as unknown, reconstructing them based on distributed measurements at $x=1$. The reformulation of the inverse problem is now as follows:
[…]
where $\mathcal{U}^{a}_{ad}$ is given by (52) and $\mathcal{K}:(a,u_{02})\in\mathcal{U}^{a}_{ad}\times\mathds{R}\mapsto\mathds{R}$ is defined as follows:
@@
We will take $\theta=1.5$, $\alpha=1$, $\beta=1$, $T=2$, $t_{1}=0$, $t_{2}=T$, $v_{0}=0$, $u_{01}=1$, $u_{02ini}=1.8$ and $aini=0.1$ as initial guesses to recover the desired value of $a_{d}=0.5$, $u_{02d}=2$ using the minimization algorithm. The numerical results can be seen in Figures 15, 15 and 16. In Figures 15 and 15, the stars and round points correspond to iterations in the computation of $u_{02}$ and $a$, respectively. With the solid line in Figure 16, we have represented the evolution of the cost. The algorithm converges well, in particular, we can see an appropriate reconstruction of the value $u_{02}$ defined on the side where we take the distributed measurements. We obtain the computed values $a_{c}=0.4999999996354925$, $u_{02c}=2.0000000002802012$ and the cost $\mathcal{K}(a_{c},u_{02c})\approx 1.e-18$.
Figure 14: Test 5, $\theta=1.5$. Iterations in the computation of $u_{02}$ by trust-region-reflective algorithm, $aini=0.1$, $u_{02ini}=1.8$.
Figure 15: Test 5, $\theta=1.5$. Iterations in the computation of $a$ in trust-region-reflective algorithm, $aini=0.1$, $u_{02ini}=1.8$.
Figure 16: Test 5, $\theta=1.5$. Evolution of the cost in trust-region-reflective algorithm, $aini=0.1$ and $u_{02ini}=1.8$.
8.4 Degeneracy power and initial data reconstruction with distributed measurements
In this sub-section, we will give a numerical simulation in line with the results 6.3 and 7.4. In particular, we assume that $u_{0}$ is of the form (54) and $v_{0}=0$, and fix the initial data $u_{01}=1$ in $(0,a)$, reconstructing the degeneracy power $\theta$ and the initial data $u_{02}$ in $(a,1)$ by distributed measurements at $x=1$. Therefore, our goal is to find $\theta\in[1,2)$ and initial data $u_{02}$ such that a solution to (51) satisfies
@@ 8.4 Degeneracy power and initial data reconstruction with distributed measurements
where $\mathcal{V}^{\theta}_{ad}=\{\theta:\theta\in[1,2-\delta]\}$ and $\mathcal{N}:(\theta,u_{02})\in\mathcal{V}^{\theta}_{ad}\times\mathds{R}\mapsto\mathds{R}$ is defined as follows:
$\mathcal{N}(\theta,u_{02})=\dfrac{1}{2}\displaystyle\int_{t_{1}}^{t_{2}}|\eta(t)-\partial_{x}u^{\theta,u_{02}}(1,t)|^{2}\,dt+\dfrac{1}{2}\displaystyle\int_{t_{1}}^{t_{2}}|\zeta(t)-\partial_{x}v^{\theta,u_{02}}(1,t)|^{2}\,dt.$
@@
We will take $a=0.5$, $\alpha=1$, $\beta=1$, $T=4$, $t_{1}=0$, $t_{2}=T$, $u_{02ini}=1.5$ and $\theta_{ini}=1.1$ as initial guesses to recover the desired value of $\theta_{d}=1.5$, $u_{02d}=2$ by the minimization algorithm. The numerical results can be seen in Figures 18, 18 and 19. The rhombus-shaped points (Figure 18) and the squares (Figure 18) correspond to iterations during the optimization algorithm in the computation of $\theta$ and $u_{02}$, respectively. With the solid line in Figure 19, we have represented the evolution of the cost. We obtain the computed value $\theta_{c}=1.5000000000000084$, $u_{02c}=2.0000000000000000$ and the cost $\mathcal{N}(\theta_{c},u_{02c})\approx 1.e-26$.
Figure 17: Test 6, $a=0.5$. Iterations in the computation of $\theta$ by trust-region-reflective algorithm, $\theta_{ini}=1.1$, $u_{02ini}=1.5$.
Figure 18: Test 6, $a=0.5$. Iterations in the computation of $u_{02}$ by trust-region-reflective algorithm, $\theta_{ini}=1.1$, $u_{02ini}=1.5$.
Figure 19: Test 6, $a=0.5$. Evolution of the cost in trust-region-reflective algorithm, $\theta_{ini}=1.1$, $u_{02ini}=1.5$.
@@
Acknowledgments: P. C. and V. D. were partially supported by the MUR Excellence Department Project awarded to the Department of Mathematics, University of Rome Tor Vergata, CUP E83C23000330006, by the PRIN 2022 PNRR-Project P20225SP98 “Some mathematical approaches to climate change and its impacts” (funded by the European Community-Next Generation EU, CUP E53D2301791 0001), and by Progetto di Ricerca di Ateneo 2024 Tor Vergata “Control and Optimization for Multilayered Artificial Neural Networks (COMANN)”, CUP E83C25000610005. V. D. was also supported by the PRIN Project 2022FPZEES “Stability in Hamiltonian Dynamics and Beyond”. P. C. was also supported by the INdAM (Istituto Nazionale di Alta Matematica) Group for Mathematical Analysis, Probability and Applications. A. D. was partially supported by Grants PID2020-114976GB-I00 and PID2024-158206NB-I00, funded by MICIU/AEI (Spain). The authors warmly thank P. Martinez and M. Yamamoto for enlightening discussions and their precious comments.